OpenStudy (anonymous):

What is the slope of the line represented by the equation y = 7?

OpenStudy (anonymous):

the equation for slope is \[m = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] and for a line y = 7 no matter what x you take the y value remains the same so you will always get \[\frac{ 7-7 }{ x _{2}-x _{1} } =\frac{ 0 }{ x _{2}-x _{1} } \] and zero divided by any number is still zero

OpenStudy (anonymous):

\[\frac{ \Delta y }{ \Delta x }\]

OpenStudy (anonymous):

What is the slope of the line represented by the equation x = -9? this is the last one

OpenStudy (anonymous):

The slope of the line is undefined...there is only one x- value for an infinite amount of y values,

OpenStudy (anonymous):

thankyou! thats the answer i thought it was, I just wanted to make sure.

OpenStudy (anonymous):

would anyone mind helping me on a few more?

OpenStudy (anonymous):

@Eleven17 No an undefined slope is the equation x = 7 because you divide by zero when calculating the slope and any number divided by zero is undefined (infinite) when zero is divided by a number the answer is still zero not undefined thus the slope can still be calculated

OpenStudy (anonymous):

@itsjustme_lol yes just close this question and post the next one as a new question

OpenStudy (anonymous):

thanks